# Eaman Jahani

### v5

@@ 4.2 Replay Analysis and Decomposition of Effects
The differences we observe in prompting behavior suggest that users are actively adapting to the capabilities of the model they are assigned. But how much of the overall performance improvement we observe for DALL-E 3 users is due to the model’s enhanced technical capacity, and how much is due to users rewriting their prompts in response to that capacity? To answer this question, we turn to the replay analysis described earlier, which allows us to isolate these two effects identified in Equation 5 empirically.
@@
Panel A of Figure 2 presents the results. To estimate the model effect, we compare the performance of prompts originally written by DALL-E 2 participants when evaluated on DALL-E 2 (the model they were written for) versus when evaluated on DALL-E 3 (Verbatim). Because these prompts were written without knowledge of DALL-E 3’s capabilities, any improvement reflects the gain from using a more capable model while holding the prompt fixed. We find that performance improves by 0.0084 in cosine similarity when these prompts are evaluated on DALL-E 3 ($p<10^{-8}$; bootstrapped standard errors clustered at the participant level), which accounts for approximately 51% of the total difference in performance between the DALL-E 2 and DALL-E 3 arms.
@@
To estimate the prompting effect, we then compare the performance of these same DALL-E 2 prompts to the performance of prompts originally written by DALL-E 3 participants, both evaluated on DALL-E 3. Because both sets of prompts are evaluated on the same model, any difference reflects the effect of users adapting their prompts to the model’s capabilities. We find that this prompting effect accounts for the remaining 48% of the total improvement, corresponding to an increase of 0.0079 in cosine similarity ($p=0.024$). The total treatment effect—i.e., the difference between the original DALL-E 2 prompts on DALL-E 2 and the original DALL-E 3 prompts on DALL-E 3—is 0.0164.
Importantly, when we apply prompts written by DALL-E 3 users to DALL-E 2, we observe no performance benefit relative to the original DALL-E 2 prompts ($\Delta=0.0020$; $p=0.56$). This asymmetry reinforces the idea that the gains from prompt adaptation depend on the model’s capacity to act on that additional information.
@@
Figure 2: Replay Analysis and Effect Decomposition. (a) Average performance of prompts evaluated across four prompt-model combinations. Comparing DALL-E 2 prompts evaluated on DALL-E 2 versus DALL-E 3 isolates the model effect. Comparing reused DALL-E 2 prompts on DALL-E 3 to original DALL-E 3 prompts on DALL-E 3 isolates the prompting effect. Error bars represent 95% confidence intervals based on bootstrapped standard errors clustered at the participant level. (b) A single target image with two submitted prompts: one written by a DALL-E 2 participant (top row) and one by a DALL-E 3 participant (bottom row). Images show how each prompt performs on both models, illustrating the model and prompting effects qualitatively.
@@
4.3 Skill Heterogeneity
@@
Total 0.0224 (0.00312) $<$ 0.00001
Total $\times$ Performance Decile -0.000115 (0.000047) 0.0152
Model 0.0113 (0.00186) $<$ 0.00001
Model $\times$ Performance Decile -0.000059 (0.000025) 0.0210
Prompting 0.0109 (0.00316) 0.000564
Prompting $\times$ Performance Decile -0.000056 (0.000048) 0.244
Table 1: Regression estimates for the total, model and prompting effects, and their interactions with participant performance decile. The total and model effects are significantly larger for lower-skilled participants, as indicated by the negative and statistically significant interaction term. The prompting effect does not vary significantly across deciles, suggesting that participants across the skill distribution adapt their prompts similarly.
Table 1 presents the results of a regression analysis that tests whether the model, prompting and total effects vary systematically across participants of differing skill levels. We regress performance on indicators for model, prompting and total effects, as well as their interactions with participants’ performance decile (ranging from lowest to highest).
We make three observations. First, the interaction between total effect and performance decile is negative and statistically significant ($-0.000115$, $p=0.0152$). This suggests that model improvements reduce the overall gap between high and low-performing users, consistent with proposition 1 in our conceptual framework. Second, the interaction between the model effect and performance decile is also negative and statistically significant ($-0.000059$, $p=0.0210$). This suggests that the model effect mainly benefits the lower-performing users—consistent with our theoretical predictions in proposition 2 and due to diminishing returns for those already near the performance ceiling. Third, we do not find any evidence that the benefits of prompt adaptation vary across the skill distribution in a meaningful way since the interaction between the prompting effect with performance decile is not statistically significant ($-0.000056$, $p=0.2444$). While this contrasts with our theoretical prediction in Proposition 2—which anticipated that higher-skill participants would benefit more from prompt adaptation—it is possible that the effect exists but is modest in size and difficult to detect given the statistical power of our design.
4.4 Prompt Revision
@@
We also evaluate whether an automated prompt adaptation system can substitute for human-driven adaptation. In the DALL-E 3 (Revised) condition, user prompts were silently rewritten by GPT-4 before being submitted to the image model. Although participants in this condition still outperformed those using DALL-E 2 ($\Delta=0.0069$; $p=0.042$), they achieved substantially lower performance than participants using DALL-E 3 without revision. On average, automated prompt revision reduced the benefit of DALL-E 3 by 58% (95% CI: [40%, 76%]). Manual inspection of the revised prompts in the DALL-E 3 (Revised) arm confirms that GPT-4 often added extraneous details or subtly altered the intended meaning of participant instructions, leading to degraded performance. In this case, the automatic rewriting process was not well-aligned with the user’s goal of replicating a specific target image as faithfully as possible. More broadly, this result highlights a potential risk of using invisible system-level prompt rewriting: when the objectives of the rewriting mechanism are not tightly aligned with user intent, even well-intentioned modifications can interfere with the user’s ability to effectively guide the model.
Together, these findings demonstrate that the performance gains observed when users are given access to a more capable generative model stem from both improvements in model architecture and changes in user behavior. Participants assigned to DALL-E 3 not only achieve higher performance but also write longer, more refined prompts over time. Our replay analysis shows that these adapted prompts account for nearly half of the overall performance improvement, with the remainder attributable to the model’s enhanced rendering capabilities. This pattern holds across the skill distribution: while lower-skilled users benefit more from model upgrades, prompt adaptation contributes consistently across all levels of baseline performance. Finally, automated prompt rewriting—when misaligned with user intent—fails to replicate the benefits of direct user adaptation. These results highlight the central role that prompting behavior plays in shaping how users interact with and benefit from improved generative AI systems.
5 Discussion and Conclusion
Our findings contribute to the literature on human–AI collaboration and the economics of technology by highlighting the role of prompt adaptation as a dynamic complement that co-evolves with model improvements (Böhm and Schedlberger 2023). As generative AI models advance—often substantially from month to month—organizations that fail to adapt their prompting strategies may forgo a meaningful share of the economic value these upgrades make possible.
@@ 5 Discussion and Conclusion
This co-evolutionary pattern is not unique to AI. It echoes dynamics observed in earlier general-purpose technologies, where technical improvements often yielded modest returns until complementary skills and practices evolved to match (Brynjolfsson 1993, David 1990). However, the pace of generative AI advancement introduces a distinctive challenge: the window for adaptation is far shorter. Organizations that treat prompting as a one-time investment rather than an ongoing capability risk failing to capture the full value of model upgrades—echoing the productivity paradox observed when complementary assets lag behind technological potential (Brynjolfsson and Hitt 2000, Brynjolfsson et al. 2021). Our findings complement this stream of work by offering direct empirical evidence that a specific dynamic complement—user adaptation, specifically through prompt refinement—accounts for roughly half of the realized performance gains associated with a model upgrade.
@@
This finding extends research on IT-enabled dynamic capabilities (Bharadwaj 2000, Joshi et al. 2010, Teece et al. 1997) and post-adoptive IT use, which emphasizes how value emerges through user experimentation and learning over time (Jasperson et al. 2005). Prompt adaptation, as we observe it, is not grounded in formal training or specialized skill. Participants in our study—none of whom were expert prompt engineers—improved performance through trial-and-error within a single session. This contrasts with earlier IT transitions, where complementary capabilities often required extensive training (Attewell 1992, Von Hippel 2006). The accessibility of prompt adaptation suggests a path toward more broadly distributed productivity gains—provided that users are supported by scaffolds and interfaces that enable iterative refinement (Rogers 2003).
At the same time, this accessibility introduces risk. Over-optimizing for a particular model version may reduce users’ ability to adapt as systems evolve. This behavioral lock-in resembles challenges in architectural innovation, where tightly coupled routines inhibit flexibility when key system components change (Henderson and Clark 1990). Supporting long-term adaptation may require not just prompt training, but workflows and learning mechanisms that encourage ongoing experimentation.
Finally, our results caution against automated prompting systems that intervene without aligning with user intent. In our experiment, DALL-E 3’s automatic prompt rewriting—designed to enhance usability—reduced realized performance gains by 58%. This does not imply that automated prompting is inherently counterproductive. In some settings, automation may simplify use or improve consistency. But when the goals of a rewriting system diverge from those of the user—as in our task, where general improvements (e.g., aesthetic quality or detail) were prioritized over faithful image replication—such interventions can interfere with the adaptive strategies users are developing. Similar challenges have been observed in other domains: for instance, in clinical documentation, automated prompt optimization improved consistency but still required manual revision to ensure quality and alignment (Yao et al. 2024). These findings reinforce the broader insight—rooted in both human–AI collaboration and dynamic capabilities research—that adaptive user behavior is not merely a response to system performance but a critical input. Interventions that disrupt this adaptive process risk eroding the very dynamic complements that make advanced systems productive (Fügener et al. 2022).
@@
While this study offers new insight into prompt adaptation, it has several limitations. First, our focus was restricted to a single transition (from DALL-E 2 to DALL-E 3) and one type of generative AI (text-to-image generation). Although our conceptual framework suggests these mechanisms may generalize to other domains, further research is needed to assess how these dynamics play out in text generation, programming, scientific research, and other high-stakes settings. Second, we observe short-run adaptation behavior in a controlled setting, whereas longer-term learning dynamics, organizational feedback structures, or team-based workflows may shape prompting strategies differently in real-world environments. Third, while our replay analysis isolates the effect of prompt adaptation, we do not identify which specific types of prompt modifications (e.g., lengthening, lexical substitution, structural rephrasing) have a causal effect on performance. Future work should investigate these finer-grained mechanisms to better inform prompting practices and interface design.
@@
Building on these limitations, several promising research directions emerge. One direction is to examine prompt adaptation longitudinally—tracking how users develop durable heuristics that generalize across domains and respond to shifting model behavior over time. Another line of inquiry involves organizational-level complements to prompting, such as shared prompt repositories, collaborative refinement practices, or analytics dashboards that surface effective patterns. A third area concerns interface design—specifically, how features like auto-complete, prompt scoring, or real-time feedback influence users’ ability to experiment and adapt. Finally, researchers might explore how organizations balance standardization (which streamlines processes) with adaptability in prompting workflows, in order to avoid creating rigid routines that lag behind technical advances.
@@
For researchers and practitioners interested in the economics of AI, our findings reinforce the idea that complements play a central role in shaping performance. The near-equal contributions of model improvements and prompt adaptation suggest that skill development and technological evolution must be treated as interdependent elements of innovation trajectories (Arthur 2009, Dosi 1982). As generative AI continues to advance, organizations that invest in adaptive, not static, complementary skills will be better positioned to realize its full value.
Appendix A Experiment Design
A.1 Task Design
Participants were asked to reproduce a single target image as closely as possible using a text-to-image generative AI model (i.e., DALL-E 2, DALL-E 3 without prompt revision, or DALL-E 3 with prompt revision, all developed by OpenAI). They did so by successively submitting prompts. In response to each submitted prompt, the model would generate an image, which was then displayed to the participant next to their assigned target image. Participants were instructed to make at least 10 attempts at trying to recreate the target image within a 25-minute window, with no upper limit on their number of attempts.
All interactions between participants and the generative AI models occurred on a custom-built online interface designed to resemble OpenAI’s ChatGPT interface but with some adjustments related to our task (e.g., displaying the target image and the total number of attempts so far to the user). On the right-hand side of the interface, participants were shown the target image they were randomly assigned to recreate. On the left-hand side, participants were shown their previously submitted prompts as well as the resulting generated images. We placed the text box where participants were able to write and submit their prompts at the bottom of the interface. Prompts were limited to a maximum of 1,000 characters. Participants were informed that their interactions with their assigned model would be memory-less, i.e., the model retained no memory of previous prompts and only used the current prompt to generate each image. Before the task, participants were provided with written and video instructions on how to interact with our experiment interface. Our task did not assume nor require prior experience with any generative AI tools.
After the task, we surveyed participants’ opinions and preferences regarding generative AI tools. We also inquired about their self-assessed occupational skills and how often they 1) engaged in creative writing, 2) wrote specific instructions, and 3) engaged in any sort of computer programming. Finally, we collected socio-demographic data, such as age, gender, and occupation.
A.2 Randomization
We randomized participants across two dimensions: the target image and the text-to-image generative AI model that participants had access to. We randomized participants across both dimensions simultaneously using complete randomization, generating 45 possible target image-model cells. We conducted a balance check after the conclusion of the experiment with a $\chi^{2}$ test across all cells. With $\chi^{2}=7.056,\;df=44$, the resulting p-value equals to 1 and thus we cannot reject the null hypothesis that the proportions are equal across all 45 groups:
$H_{0}:p_{1}=p_{2}=\cdots=p_{45}$
Participants were unaware of this randomization.
A.2.1 Generative Models
We randomly assigned participants to 1 of 3 generative models:

### v6

@@ 4.1.2 Replay Analysis and Decomposition of Effects
The differences we observe in prompting behavior suggest that users are actively adapting to the capabilities of the model they are assigned. But how much of the overall performance improvement we observe for DALL-E 3 users is due to the model’s enhanced technical capacity, and how much is due to users rewriting their prompts in response to that capacity? To answer this question, we turn to the replay analysis described earlier, which allows us to isolate these two effects empirically.
@@
Panel A of Figure 2 presents the results. To estimate the model effect, we hold prompts fixed and compare how the same DALL-E 2 prompts perform when evaluated on DALL-E 2 versus when replayed on DALL-E 3 (Verbatim). Because these prompts were written without knowledge of DALL-E 3’s capabilities, any improvement reflects the gain from using a more capable model while holding the prompt fixed. We find that performance improves by 0.0084 in cosine similarity when these prompts are evaluated on DALL-E 3 ($p<10^{-8}$; bootstrapped standard errors clustered at the participant level), which accounts for approximately 51% of the total difference in performance between the DALL-E 2 and DALL-E 3 arms.
@@
To estimate the prompting effect, we then compare the performance of these same DALL-E 2 prompts to the performance of prompts originally written by DALL-E 3 participants, both evaluated on DALL-E 3. Because both sets of prompts are evaluated on the same model, any difference reflects the effect of users adapting their prompts to the model’s capabilities. We find that this prompting effect accounts for the remaining $\approx 48$% of the total improvement [^10] The total, model, and prompting effects are estimated using three separate two-way fixed-effects models, as explained in Appendix F.1.3. Because each effect is estimated from a different subset of counterfactual prompt–model comparisons with its own fixed-effects, the resulting model and prompting effect estimates are not guaranteed to sum exactly to the total improvement., corresponding to an increase of 0.0079 in cosine similarity ($p=0.024$). Importantly, when we apply prompts written by DALL-E 3 users to DALL-E 2, we observe no performance benefit relative to the original DALL-E 2 prompts ($\Delta=0.0020$; $p=0.56$). This asymmetry reinforces the idea that the gains from prompt adaptation depend on the model’s capacity to act on that additional information.
Figure 2: Replay Analysis and Effect Decomposition for the Image Replication Task.
@@
(a) Average performance of prompts evaluated across four prompt-model combinations. Comparing DALL-E 2 prompts evaluated on DALL-E 2 versus DALL-E 3 isolates the model effect. Comparing reused DALL-E 2 prompts on DALL-E 3 to original DALL-E 3 prompts on DALL-E 3 isolates the prompting effect. Error bars represent 95% confidence intervals based on bootstrapped standard errors clustered at the participant level. (b) A single target image with two submitted prompts: one written by a DALL-E 2 participant (top row) and one by a DALL-E 3 participant (bottom row). Images show how each prompt performs on both models, illustrating the model and prompting effects qualitatively.
@@
4.1.3 Distributional Effects
Table 1: Distributional Effects of Model Upgrades in the Image Replication Task
@@ 4.1.3 Distributional Effects
Total 0.022 (0.003) << $10^{-5}$
Total $\times$ Outcome Quantile (Percentile) -0.00012 (0.00005) 0.015
Model 0.011 (0.002) << $10^{-5}$
Model $\times$ Outcome Quantile (Percentile) -0.00006 (0.00003) 0.021
Prompting 0.011 (0.003) << $10^{-5}$
Prompting $\times$ Outcome Quantile (Percentile) -0.00006 (0.00005) 0.244
Negative interactions with outcome quantile indicate that gains from model upgrades are larger at lower points in the realized outcome distribution, implying compression of performance outcomes. The total and model effects exhibit statistically significant compression, while we do not detect a statistically significant distributional pattern for the prompting effect. Note: the un-interacted total effect coefficient gives the QTE estimate at the 0th percentile.
Next, we examine how model improvements and prompt adaptation reshape the distribution of performance outcomes in the image replication task using a quantile treatment effects (QTE) perspective.
Table 1 reports estimates of the total, model, and prompting effects at different points of the realized outcome distribution, with outcome quantiles indexed by percentile. The total effect exhibits a negative and statistically significant interaction with outcome quantile (0.000115-0.000115, $p=0.0152$), indicating that performance gains from upgrading from DALL-E 2 to DALL-E 3 are larger at lower points in the outcome distribution. This pattern is consistent with the framework’s prediction that model improvements compress outcomes in bounded tasks. This compression is driven primarily by the model effect. Holding prompts fixed, the interaction between the model effect and outcome quantile is also negative and statistically significant (0.000059-0.000059, $p=0.0210$), consistent with diminishing returns near the performance ceiling. While the bounded formulation predicts that prompt adaptation exerts an expanding force on the outcome distribution, we do not detect a statistically significant interaction between the prompting effect and outcome quantile in this setting (0.000056-0.000056, $p=0.2444$). This result does not rule out the presence of an expanding prompting channel, but implies that any such effect is modest in magnitude relative to the dominant equalizing model effect.
4.1.4 Prompt Revision
@@
Finally, we examine whether automated prompt rewriting can serve as a substitute for human-driven prompt adaptation in the image replication task. In the DALL-E 3 (Revised) condition, user prompts were silently rewritten by GPT-4 before being submitted to the image model. Although participants in this condition still outperformed those using DALL-E 2 ($\Delta=0.0069$; $p=0.042$), they achieved substantially lower performance than participants using DALL-E 3 without rewriting. On average, automated prompt revision reduced the benefit of DALL-E 3 by 58% (95% CI: [40%, 76%]). Qualitative inspection of revised prompts suggests that the rewriting process often introduced additional details or altered phrasing in ways that were misaligned with participants’ objective of replicating a specific target image as faithfully as possible. In this bounded, steerable task, such modifications reduced users’ ability to precisely control the model’s output and partially offset the performance gains achieved through direct user-side prompt adaptation.
4.2 Logo Generation Task
We next turn to the logo generation task. This logo generation task, however, represents an open-ended, creative application of generative AI—such as design or ideation—in which success is evaluated comparatively rather than against a known target and output quality does not have a natural upper bound. Our empirical analyses mirror that of the image replication experiment.
Figure 3: Overall Performance and Prompting Behavior for the Logo Generation Task
(A) For three example target images, the middle row shows participant-generated images closest to the mean Bradley-Terry $\beta$ across all prompts. The rows above and below show images of approximately one average treatment effect (ATE) higher or lower quality, illustrating the typical performance difference between model conditions. (B) Top: average Bradley-Terry $\beta$, separately for DALL-E 2 and DALL-E 3 participants. Bottom: the difference between these averages, with the dark blue line indicating the overall ATE and the shaded region showing the 95% confidence interval. (C) Average prompt length by attempt (y-axis), with error bars representing 95% confidence intervals. Color shading indicates the average textual similarity between each prompt and the participant’s previous prompt, capturing the extent of prompt reuse and refinement over time.
4.2.1 Overall Impact of Model Upgrades
We begin by examining whether access to a more capable model improves performance in the logo generation task. Participants assigned to DALL-E 3 produce substantially higher-quality logos than those assigned to DALL-E 2, with an estimated average treatment effect of 0.8075 in Bradley–Terry $\beta$ (95% CI: [0.7120, 0.9030], $p<10^{-10}$), corresponding to an improvement of approximately 0.57 standard deviations. Panel A of Figure 3 provides qualitative intuition for the magnitude of this effect by showing representative logos at the mean, as well as approximately one average treatment effect above and below the mean. Panel B of Figure 3 shows how this performance gap unfolds over successive prompting attempts. As in the image replication task, the advantage of the more capable model appears early and persists, although it does not appear to diverge across attempts. Finally, Panel C of Figure 3 shows similar differences to the image replication task in prompting behavior across model conditions, with participants using DALL-E 3 prompts that are approximately 24.7% longer on average and exhibiting greater iteration-to-iteration similarity, consistent with more exploitative refinement. The proportion of nouns and adjectives is also once again statistically indistinguishable across model conditions (43% for DALL-E 3 vs. 42% for DALL-E 2; $p$ = 0.12)
Together, these patterns show that in an open-ended, creative task without a fixed target or performance ceiling, increased model capability leads to large and persistent performance gains. These gains are accompanied by systematic changes in user prompting behavior as users adapt their prompts in response to the model’s perceived capabilities. These results largely mirror those from the image replication task.
4.2.2 Replay Analysis and Decomposition of Effects
Next, we apply the same replay-based decomposition used in the image replication task to separate the overall performance gains in the logo generation experiment into components attributable to model improvements and to user prompt adaptation. Panel A of Figure 4 presents the resulting decomposition.
The overall average treatment effect of upgrading from DALL-E 2 to DALL-E 3 is $0.8075$ in Bradley–Terry $\beta$. In contrast to the image replication task, the vast majority of this gain is attributable to the model effect: replaying prompts written by DALL-E 2 participants on DALL-E 3 increases performance by $0.748$ ($p<10^{-10}$), accounting for approximately $92.7\%$ of the total effect. The corresponding prompting effect is small in magnitude ($0.0577$), statistically indistinguishable from zero ($p=0.101$), and accounts for only $7.2\%$ of the total effect. Consistent with this pattern, replaying prompts written for DALL-E 3 on DALL-E 2 yields no performance improvement ($\Delta\beta=-0.0125$, $p=0.62$).
Panel B of Figure 4 provides a qualitative illustration of this pattern for a representative logo brief. Although the prompt written by the DALL-E 3 participant is longer and more detailed—consistent with the prompt adaptation patterns shown in Panel C of Figure 3—it does not perform substantially better on DALL-E 3 than the prompt written by the DALL-E 2 participant. Instead, the primary performance difference arises from evaluating either prompt on the more capable model. While this comparison is illustrative rather than representative, it highlights how prompt adaptation in this task does not translate into large additional gains beyond those delivered by the model itself.
Taken together, these results indicate that in the logo generation task, realized performance gains are driven primarily by improvements in model capability, with prompt adaptation playing a comparatively limited role.
Figure 4: Replay Analysis and Effect Decomposition for the Logo Generation Task.
(a) Average performance of prompts evaluated across four prompt-model combinations. Comparing DALL-E 2 prompts evaluated on DALL-E 2 versus DALL-E 3 isolates the model effect. Comparing reused DALL-E 2 prompts on DALL-E 3 to original DALL-E 3 prompts on DALL-E 3 isolates the prompting effect. Error bars represent 95% confidence intervals based on bootstrapped standard errors clustered at the participant level. (b) A single client description with two submitted prompts: one written by a DALL-E 2 participant (top row) and one by a DALL-E 3 participant (bottom row). Images show how each prompt performs on both models, illustrating the model and prompting effects qualitatively.
4.2.3 Distributional Effects
Table 2: Distributional Effects of Model Upgrades in the Logo Generation Task
Effect Estimate (SE) p-value
Total 1.135 (0.050) << $10^{-5}$
Total $\times$ Outcome Quantile (Percentile) -0.0065 (0.0008) << $10^{-5}$
Model 1.012 (0.035) << $10^{-5}$
Model $\times$ Outcome Quantile (Percentile) -0.0053 (0.0005) << $10^{-5}$
Prompting 0.122 (0.050) 0.014
Prompting $\times$ Outcome Quantile (Percentile) -0.0013 (0.0008) 0.117
Negative interactions with outcome quantile indicate that gains from model upgrades are larger at lower points in the realized outcome distribution, implying compression of performance outcomes. As in the image replication task, distributional compression is driven by the model effect, while we do not detect a statistically significant distributional pattern for the prompting effect. Note: the un-interacted total effect coefficient gives the QTE estimate at the 0th percentile.
We also apply the same QTE approach used in the image replication task to examine how the model upgrade reshapes the distribution of performance outcomes in the logo generation experiment. Table 2 reports estimates of the total, model, and prompting effects at different points of the realized outcome distribution, with outcome quantiles indexed by percentile.
As was the case in the image replication task, the total effect of upgrading from DALL-E 2 to DALL-E 3 exhibits a negative and statistically significant interaction with outcome quantile, indicating that performance gains are larger at lower points in the outcome distribution and that the upgrade compresses the distribution of outcomes. This pattern is again driven by the model effect: holding prompts fixed, the interaction between the model effect and outcome quantile is negative and statistically significant. By contrast, we do not detect a statistically significant interaction between the prompting effect and outcome quantile. Although this lack of statistical significance does not rule out the compressive impact of prompt adaptation predicted by the unbounded formulation of the model, it places an upper bound on the magnitude of any such effect in this setting.
Taken together, these results indicate that, as in the image replication task, distributional compression in the logo generation experiment arises primarily from improvements in model capability, with prompt adaptation playing a comparatively limited role in shaping the distribution of outcomes.
4.2.4 Prompt Revision
We conclude analyses of this experiment by examining the effects of automated prompt rewriting in the logo generation task. As in the image replication task, participants assigned to DALL-E 3 with automatic prompt revision outperform those using DALL-E 2, with an average treatment effect of $0.8337$ in BT $\beta$ ($p<10^{-10}$). However, in contrast to the image replication task, automated prompt rewriting does not reduce performance relative to DALL-E 3 without rewriting; instead, it is associated with a $3.2\%$ increase in BT $\beta$ on average (95% CI: $5.6\%,12.1\%]$).
One possible interpretation is that in an open-ended, creative task, the specific prompt rewriting used here, which primarily embellishes and elaborates user inputs, can enrich prompts in ways that facilitate exploration, introduce useful variation, or highlight design elements that improve perceived quality. While we do not directly observe these mechanisms, the positive effect of rewriting in this task is consistent with the idea that automated prompt modification can complement user prompting when it is aligned with users’ goals and the structure of the task.
5 Discussion
In this paper, we show that realized performance gains from generative AI upgrades reflect an interaction between model capability, user adaptation, and task structure. In a bounded, steerable task with an objective target (image replication), improvements in model quality translated into performance gains through two channels: roughly half of the improvement arose mechanically from the stronger model, while the other half reflected users adapting their prompts to better exploit the model’s capabilities. In this setting, prompt adaptation is a substantive and economically meaningful complement to model improvement. By contrast, in an open-ended creative task without a natural performance ceiling (logo generation), performance gains were driven primarily by improvements in model capability. Although users again modified their prompting behavior when interacting with the stronger model, replay analyses indicate that prompt adaptation accounted for only a small share of the overall improvement. Thus, we argue that prompt adaptation functions as a dynamic complement to generative AI models, with task structure emerging as a key boundary condition: it plays a first-order role when success is defined against a fixed objective, but a more limited role when quality is open-ended and comparative.
A related question we address is whether prompt adaptation must occur through direct user experimentation or whether it can be partially automated by the system itself. In practice, many generative AI systems implement automated prompt rewriting to reduce user effort, implicitly positioning automation as a substitute for user-driven adaptation. Our findings suggest that the effectiveness of such automation depends critically on task structure and alignment with user goals. In the bounded image replication task, the DALL-E 3 endpoint’s automated prompt rewriting reduced realized gains, consistent with misalignment between the rewriting logic and users’ objective of faithfully reproducing a target image. By contrast, in the open-ended logo generation task, the same rewriting mechanism modestly improved performance. Together, these findings underscore the importance of careful design in automated prompt rewriting systems. When system prompts are well aligned with user goals and task structure, automated rewriting can improve performance. but poorly aligned system prompts, especially when hidden, can degrade performance and/or limit users’ ability to discover or exploit emergent but useful ways of interacting with a generative AI system. Such complexities are highlighted by related work, such as yao2024clinicians, which studies automated prompt optimization in the context of clinical documentation.
@@ 5 Discussion
More generally, our results both align with and help clarify themes in the literature on human–AI collaboration and the economics of technology, particularly the role of dynamic complements such as prompt adaptation that co-evolve with technical change (bohm2023co-evolution). As generative AI models advance, often substantially from month to month, organizations that fail to adapt their prompting strategies in task settings where precision and steerability matter may forgo a meaningful share of the economic value these upgrades make possible. This co-evolutionary pattern is not unique to AI. It echoes dynamics observed in earlier general-purpose technologies, where technical improvements often yielded modest returns until complementary skills and practices evolved to match (brynjolfsson1993productivity, david1990dynamo). However, the pace of generative AI advancement introduces a distinctive challenge: the window for adaptation is far shorter. Organizations that treat prompting as a one-time investment rather than an ongoing capability risk failing to capture the full value of model upgrades—echoing the productivity paradox observed when complementary assets lag behind technological potential (brynjolfsson2000beyond, brynjolfsson2021productivity). Our findings complement this stream of work by offering direct empirical evidence that a specific dynamic complement—user adaptation, specifically through prompt refinement—can account for a substantial share of realized performance gains from model upgrades in task settings where precision and steerability matter.
@@
Seen from a broader organizational perspective, these findings also speak to long-standing work on IT-enabled dynamic capabilities (bharadwaj2000resource, joshi2010changing, teece1997dynamic) and post-adoptive IT use (jasperson2005comprehensive). Prompt adaptation, as we observe it, is not grounded in formal training or specialized skill. Participants in our study, drawn from a general population rather than a specialized group of prompt engineers, improved performance through trial-and-error within a single session. This contrasts with earlier IT transitions, where complementary capabilities often required extensive training (attewell1992technology, von2006democratizing). The relative accessibility of prompt adaptation highlights one potential pathway toward more broadly distributed productivity gains, provided that users are supported by scaffolds and interfaces that enable iterative refinement (rogers2003diffusion). At the same time, this accessibility introduces risk. Over-optimizing for a particular model version may reduce users’ ability to adapt as systems evolve. This behavioral lock-in resembles challenges in architectural innovation, where tightly coupled routines inhibit flexibility when key system components change (henderson1990architectural). Supporting long-term adaptation may require not just prompt training, but workflows and learning mechanisms that encourage ongoing experimentation.
@@
Our study has several limitations. First, our experimental interface was intentionally memoryless: each prompt was processed independently, and participants were informed that prior interactions would not influence subsequent generations. This simplification was necessary to isolate prompt-level adaptation in a controlled setting, but it departs from the conversational context that characterizes many real-world AI systems; our estimates therefore characterize performance and adaptation conditional on a memoryless interaction structure. Second, our focus was restricted to a single transition (from DALL-E 2 to DALL-E 3) and one type of generative AI (text-to-image generation). Although our conceptual framework suggests these mechanisms may generalize to other domains, further research is needed to assess how these dynamics play out in text generation, programming, scientific research, and other high-stakes settings. Third, we observe short-run adaptation behavior in a controlled setting, whereas longer-term learning dynamics, organizational feedback structures, or team-based workflows may shape prompting strategies differently in real-world environments. Finally, while our replay analysis helps isolate the contribution of prompt adaptation relative to model improvements, it does not fully disentangle the causal effects of specific prompt modifications (e.g., lengthening, lexical substitution, structural rephrasing), particularly given the path-dependent nature of user adaptation.
@@
Building on these limitations, several promising research directions emerge. One important next step is to examine prompt adaptation in conversational, memory-enabled settings, where prior interactions persist and can be leveraged for cumulative learning by both users and models, and to study how this co-evolution impacts performance. Relatedly, future studies might explore prompt adaptation longitudinally, observing how users develop durable heuristics that generalize across domains and respond to shifting model behavior over time. Another line of inquiry involves organizational-level complements to prompting, such as shared prompt repositories, collaborative refinement practices, or analytics dashboards that surface effective patterns. A third area concerns interface design—specifically, how features like auto-complete, prompt scoring, or real-time feedback influence users’ ability to experiment and adapt. Finally, researchers might explore how organizations balance standardization (which streamlines processes) with adaptability in prompting workflows, in order to avoid creating rigid routines that lag behind technical advances.
@@
For researchers and practitioners interested in the economics of AI, our findings reinforce the idea that complements play a central role in shaping performance. In some task settings, prompt adaptation can account for as much as half of the realized performance gains from model upgrades, underscoring that skill development and technological evolution must be treated as interdependent elements of innovation trajectories (arthur2009nature, dosi1982technological). As generative AI continues to advance, organizations that invest in adaptive, not static, complementary skills will be better positioned to realize its full value.
Appendix A Conceptual Framework and Derivations
This appendix develops the formal derivations and proofs underlying the conceptual framework presented in the main text. Whereas the main text emphasizes intuition and empirical implications, the goal here is to make the underlying mechanics transparent.
We first analyze a bounded quality function, where output quality is capped at 1, and then extend the analysis to unbounded tasks. The exposition below focuses on the bounded case; the unbounded case follows by analogy.
A.1 Model Setup
Consider a task where users attempt to replicate a target object (e.g., an image) using a generative model. Perfect replication yields a quality of 1. Define the quality function
$Q(\theta,s,x)=1-\exp[-\theta sx],$
where $\theta\in(0,1]$ denotes model capacity; $s\in(0,1]$ denotes user skill at writing prompts; and $x\geq 0$ denotes the user’s effort in generating prompts.
This specification has two desirable properties: